# Learning Heuristics for Transit Network Design and Improvement with Deep Reinforcement Learning

### v1

@@ 6.3 Results
In addition, NEA’s balanced-case networks reduce the average passenger trip time by 34% versus the STL network at the same time as reducing costs. They also reduce the number of transfers that passengers have to make, especially reducing the percentage of trips that require three or more transfers from 41.83% to 36.41%. Trips requiring three or more transfers are widely regarded as being so unattractive that riders will not make them (but note that these trips are still included in the calculation of $C_{p}$). So NEA’s networks may increase overall transit ridership. This not only makes the system more useful to the city, but also increases the agency’s revenue from fares.
These results demonstrate that learned heuristics may be useful in complex real-world transit planning scenarios. They can improve on an existing transit system in multiple dimensions, in a way sensitive to the planner’s preference over these dimensions. Transit agencies may be able to use it to “do more with less”, offering better service at reduced cost.
@@
We also note that PC-EA achieves lower average trip time $C_{p}$ than both STL and NEA for all three $\alpha$ values, but at the cost of worsening total route time $C_{o}$ versus STL in each case. Furthermore, at $\alpha=1.0$ where we disregard $C_{o}$, it achieves only 3% lower $C_{p}$ than NEA. This suggests that while the path-combining heuristic for assembling routes is indeed helpful on its own, the learned heuristics of the neural net policy contribute significantly to its performance, reinforcing our observations in subsubsection 5.2.3 and subsection 5.3.
@@
In these experiments, the nodes of the graph represent census dissemination areas instead of existing bus stops. So in order to be used in reality, the proposed routes would need to be fitted to the existing locations of bus stops in the dissemination areas, likely making multiple stops within each area. To minimize the costs of the network redesign, it would be necessary to keep stops at most or all stop locations that have shelters, as constructing or moving these may cost tens of thousands of dollars, while stops with only a signpost at their locations can be moved at the cost of only a few thousand dollars. An algorithm for translating our census-level routes to real-stop-level routes, in a way that minimizes cost, is an important next step for making this work applicable. We leave this for future work.
@@ 7 CONCLUSIONS
The choice of low-level heuristics has a major impact on the effectiveness of a metaheuristic algorithm, and our results demonstrate that using deep reinforcement learning to learn heuristics can offer substantial improvements over, or alongside with, engineering them directly. They also demonstrate that learned heuristics may be better-suited to optimizing the transit networks of real-world cities than are human-engineered heuristics, potentially allowing transit agencies to reduce costs while improving quality of service. The method we propose here is a simple one, and presents several obvious directions for improvement.
@@
One is that the construction process on which our neural net policy is trained differs from the metaheuristic algorithm (an improvement process) in which it is deployed. Another is that we learn just one type of heuristic, that for constructing a route from shortest-path segments, though we note that the neural net may be learning multiple distinct heuristics for how to do this under different circumstances. A natural next step would be to train a more diverse set of neural heuristics, perhaps including route-lengthening and -shortening operators, and to do so directly in the context of an improvement process, which may make the learned heuristics more appropriate to that process.
@@
A remaining question is why the random path-combining heuristic used in the PC-EA experiments outperforms the learned heuristic in the extreme of the passenger perspective. As noted in subsubsection 5.2.3, in principle the neural net should be able to learn whatever policy the random path-combiner is enacting: in the limit, it could learn to give the same probability to all actions when $\alpha=1.0$. We wish to explore in more depth why this does not occur, and whether changes to the policy architecture or learning algorithm might improve the learned heuristic’s performance to match or exceed the random policy’s.
@@
The evolutionary algorithm in which we use our learned heuristic was chosen for its speed and simplicity, which aided in implementation and rapid experimentation. But as noted in subsection 4.2, other more costly but state-of-the-art algorithms, like NSGA-II, may allow still better performance to be gained from learned heuristics. Future work should attempt to use learned heuristics like ours in more varied metaheuristic algorithms.
@@
Another promising place where machine learning can be applied is to learn hyper-heuristics for a metaheuristic algorithm. A typical hyper-heuristic, as used by Ahmed et al. (2019) and Hüsselmann et al. (2023), adapts the probability of using each low-level heuristic based on its performance, while the algorithm is running. Instead of the simple methods used in existing work, we would like to train a neural net policy to act as a hyper-heuristic to select among low-level heuristics, given details about the scenario and previous steps taken during the algorithm.
@@
Our aim in this paper was to show whether learned heuristics can be used in a lightweight metaheuristic algorithm to improve its performance. Our results show that they can, and in fact that they give performance competitive with, and in some cases better than, state-of-the-art algorithms using human-engineered heuristics. Our results provide compelling evidence that they may help transit agencies substantially reduce operating costs while delivering better service to riders in real-world scenarios.
@@ Appendix A Neural Net Architecture
In this section we describe the architecture of the neural net used in our Learned Constructor policy $\pi_{\theta}$.
@@ A.1 Input Features
The policy network operates on three types of input: an $n\times 4$ matrix of node feature vectors $X$, an $n\times n\times 13$ tensor of edge features $E$, and a global state vector $\mathbf{s}$.
@@
An edge feature $e_{ij}$ is composed of:
@@
  • $\textup{self}=(i=j)$, a binary feature indicating whether this edge connects a node to itself
@@
  • $\tau_{\mathcal{R}ij}$ if $c_{0ij}=1$, 00 otherwise
  • $\tau_{ij}$ if $s_{ij}=1$, 00 otherwise
  • $D_{ij}$, the demand between the nodes
@@ A.2 Policy Network Architecture
The policy $\pi_{\theta}$ is a neural net with three components: a GNN “backbone”, a halting module $\textup{NN}_{halt}$, and an extension module $\textup{NN}_{ext}$. The halting and extension modules both operate on the outputs from the GNN. All nonlinearities are ReLU functions. All attention modules are multi-headed with 4 heads. A common embedding dimension of $d_{embed}=64$ is used; unless otherwise specified, each component of the system outputs vectors of this dimension.
@@ A.2.1 Backbone network
The backbone is a graph attention net composed of five GATv2(Brody et al., 2021) layers separated by ReLU nonlinearities. The network takes as input the node feature collection $X$ and the edge features $E$, and the final layer outputs a collection of node embeddings $Y=\{\mathbf{y}_{i}\;\forall i\in\mathcal{N}\}$. These node embeddings are used by the two policy heads to compute action probabilities.
@@ A.2.2 Halting module
$\textup{NN}_{halt}$ is a simple feed-forward MLP. It takes as input the concatenation of the node embeddings of first and last nodes on the current route, the mean of the node embeddings over the graph, $\mathbf{s}_{t}$ and $\tau_{r_{t}}$. It outputs a scalar $h$. We apply the sigmoid function to $h$ get the halting policy:
$\pi(\textup{halt})=\sigma(h)=\frac{1}{1+e^{-h}},\;\pi(\textup{continue})=1-\sigma(h)$ (10)
@@ A.2.3 Extension module
The extension module $\textup{NN}_{ext}$ is finely tuned to the structure of the cost function. Consider first the case where $\alpha=1$. We observe that the quality of any candidate route $r$ is then solely a function of the edges $(i,j,\tau_{rij})$ that it adds to the set $\mathcal{E}_{\mathcal{R}}$ of direct-transit-connection edges, where $\tau_{rij}$ denotes the time to get from $i$ to $j$ on route $r$. Therefore, for each path $a\in\textup{SP}$ that we consider as an extension of $r$, a reasonable heuristic is to compute a scalar score $o_{ij}$ for each edge $(i,j,\tau_{aij})$, and let the path’s score $o_{a}$ be the sum of the scores of all pairs of nodes that are connected by that path.
Following this intuition, for each node pair $(i,j)$, we concatenate $[\mathbf{y}_{i},\mathbf{y}_{j},\mathbf{e}_{ij},\tau{(r+a)ij}]$ into a vector and apply an MLP to to compute the scalar $o_{ij}$. For each path $a\in\textup{SP}$, we sum these node-pair scores to get path scores: $o_{a}=\sum_{i,j\in a}o_{ij}$.
If $\alpha<1$, then the quality of $r$ still depends on the edges it adds to $\mathcal{E}_{\mathcal{R}}$, but also on its time, and whether it is a good choice to extend $r$ by $a$ depends generally on the state $\mathcal{S}$. So, we concatenate $[o_{a},\tau_{a},\mathbf{s}_{t}]$ and apply another MLP to obtain a final score $\hat{o}_{a}$ for each candidate path. The extension policy is then the softmax of these values:
@@ A.4 Training Hyper-Parameters
Training was performed using the well-known Adam optimizer Kingma and Ba (2015). Table 7 gives the hyper-parameters used during training. In each epoch of training, we augment the training dataset by applying a set of random transformations. These include rescaling the node positions and by a uniformly-sampled random factor uniformly sampled in the range $[0.4,1.6]$, rescaling the demand magnitudes by a random factor uniformly sampled in the range $[0.8,1.2]$, mirroring node positions about the $y$ axis with probability $0.5$, and rotating the node positions by a random angle in $[0,2\pi)$ about their geometric center.
We also randomly sample $\alpha$ separately for each training city in each epoch. Each time alpha is sampled, it has equal probability of being set to $0.0$, set to $1.0$, and sampled uniformly in the range $[0,1]$. This is to encourage the policy to learn both to handle the extreme cases of the passenger perspective and operator perspective, as well as intermediate cases.

### v2

@@ 6.3.1 Results
In addition, NEA’s balanced-case networks reduce the average passenger trip time by 34% versus the STL network, and reduce the number of transfers that passengers have to make, especially reducing the percentage of trips that require three or more transfers from 41.83% to 36.41%. Trips requiring three or more transfers are widely regarded as being so unattractive that riders will not make them (but note that these trips are still included in the calculation of $C_{p}$). So NEA’s networks may even increase overall transit ridership. This not only makes the system more useful to the city’s residents, but also increases the agency’s revenue from fares.
@@
We also note that PC-EA achieves lower average trip time $C_{p}$ than both STL and NEA for all three $\alpha$ values, but at the cost of worsening total route time $C_{o}$ versus STL in each case. Furthermore, at $\alpha=1.0$ where we disregard $C_{o}$, it achieves only 3% lower $C_{p}$ than NEA. This supports our conclusions in subsection 5.2.3 and subsection 5.3 that the learned heuristics of the neural net policies $\pi_{\theta}$ are the main driver of NEA’s performance, rather than simply the process of assembling routes from shortest paths.
@@
In these experiments, the nodes of the graph represent census dissemination areas instead of existing bus stops. In order to be used in reality, the proposed routes would need to be fitted to the existing locations of bus stops in the CDAs, likely making multiple stops within each CDAs. To minimize the cost of the network redesign, it would be necessary to keep stops at most or all stop locations that have shelters, as constructing or moving these may cost tens of thousands of dollars, while stops with only a signpost at their locations can be moved at the cost of only a few thousand dollars. An algorithm for translating our census-level routes to real-stop-level routes, in a way that minimizes cost, is an important next step for making this work applicable. We leave this for future work.
@@ 7 CONCLUSIONS
The choice of low-level heuristics has a major impact on the effectiveness of a metaheuristic algorithm, and our results show that using deep reinforcement learning to learn heuristics can offer substantial benefits when used alongside human-engineered heuristics. They also show that learned heuristics may be useful in complex real-world transit planning scenarios. They can improve on an existing transit system in multiple dimensions, in a way sensitive to the planner’s preference over these dimensions, and so may enable transit agencies to offer better service at reduced cost.
@@
Our heuristic-learning method has several limitations. One is that the construction MDP on which our neural net policies are trained differs from the evolutionary algorithm - an improvement process - in which it is deployed. Another is that it learns just one type of heuristic, that for constructing a route from shortest paths, though we note that the neural net may be learning multiple distinct heuristics for how to do this under different circumstances. Natural next steps would be to learn a more diverse set of heuristics, perhaps including route-lengthening and -shortening operators, and to train our neural net policies directly in the context of an improvement process, which may yield learned heuristics that are better-suited to that process.
@@
A remaining question is why the random path-combining heuristic used in the PC-EA experiments outperforms the learned heuristic in the extreme of the passenger perspective. As noted in subsection 5.2.3, in principle the neural net should be able to learn whatever policy the random path-combiner is enacting: in the limit, it could learn to give the same probability to all actions when $\alpha=1.0$. We wish to explore in more depth why this does not occur, and whether changes to the policy architecture or learning algorithm might allow the learned heuristic’s performance to match or exceed that of the random policy.
@@
The evolutionary algorithm in which we use our learned heuristic was chosen for its speed and simplicity, which aided in implementation and rapid experimentation. But as noted in subsection 4.2, other more costly state-of-the-art algorithms, like NSGA-II, may allow still better performance to be gained from learned heuristics. Future work should attempt to use learned heuristics like ours in more varied metaheuristic algorithms.
@@
Another promising place where machine learning can be applied is to learn hyper-heuristics for a metaheuristic algorithm. A typical hyper-heuristic, as used by Ahmed et al. (2019) and Hüsselmann et al. (2023), adapts the probability of using each low-level heuristic based on its performance, while the algorithm is running. It would be interesting to train a neural net policy to act as a hyper-heuristic to select among low-level heuristics, given details about the scenario and previous steps taken during the algorithm.
@@
Our aim in this paper was to show whether learned heuristics can be used in a lightweight metaheuristic algorithm to improve its performance. Our results show that they can, and in fact that they give performance competitive with, and in some cases better than, state-of-the-art algorithms using human-engineered heuristics. This is compelling evidence that they may help transit agencies substantially reduce operating costs while delivering better service to riders in real-world scenarios.
@@ Appendix A Neural Net Architecture
In this appendix, we give a detailed description of the architecture of the neural net used in our GNN policy $\pi_{\theta}$.
@@ A.1 Input Features
The policy net operates on three inputs: an $n\times 4$ matrix of node feature vectors $X$, an $n\times n\times 13$ tensor of edge features $E$, and a global state vector $\mathbf{s}$.
@@
An edge feature $\mathbf{e}_{ij}$ is composed of the following elements:
  • $D_{ij}$, the demand between the nodes
@@
  • $\textup{self}=1$ if $i=j$, 00 otherwise
@@
  • $\tau_{rij}$ if $c_{0ij}=1$ where $r$ the route that provides the shortest direct trip between $i$ and $j$, 00 otherwise
@@
Note that $\alpha$ and $1-\alpha$ are included in both $\mathbf{s}$ and $\mathbf{e}_{ij}$.
@@ A.2 Policy Network Architecture
The policy $\pi_{\theta}$ is a neural net with three components: a GAT “backbone”, a halting module $\textup{NN}_{halt}$, and an extension module $\textup{NN}_{ext}$. The halting and extension modules both operate on the outputs from the GNN. All nonlinearities are ReLU functions. A common embedding dimension of $d_{\text{embed}}=64$ is used throughout; unless otherwise specified, each component of the system outputs vectors of this dimension.
@@ A.2.1 Backbone network
The backbone is a graph attention net composed of five GATv2(Brody et al., 2021) multi-head graph attention layers separated by ReLU nonlinearities. 4 heads are used in each multi-head attention layer. The network takes as input the node feature collection $X$ and the edge features $E$, and the final layer outputs a collection of node embeddings $Y=\{\mathbf{y}_{i}\;\forall i\in\mathcal{N}\}$. These node embeddings are used by the two policy heads to compute action probabilities.
@@ A.2.2 Halting module
$\textup{NN}_{halt}$ is a simple feed-forward MLP with 2 hidden layers. It takes as input the concatenation of the node embeddings of the first and last nodes on the current route, $\mathbf{s}_{t}$ and $\tau_{r_{t}}$. It outputs a scalar $h$. We apply the sigmoid function to $h$ get the halting policy:
$\pi(\textup{halt}|s_{t})$ $=\sigma(h)=\frac{1}{1+e^{-h}}$ (11)
$\pi(\textup{continue}|s_{t})$ $=1-\sigma(h)$ (12)
@@ A.2.3 Extension module
The extension module $\textup{NN}_{ext}$ is finely tuned to the structure of the cost function. Consider first the case where $\alpha=1$. The benefit of a candidate route $r$ is then solely a function of the edges $(i,j,\tau_{rij})$ that it adds to the set $\mathcal{E}_{\mathcal{R}}$ of direct-transit-connection edges, where $\tau_{rij}$ denotes the time to get from $i$ to $j$ on route $r$. Therefore, for each path $a\in\textup{SP}$ that we consider as an extension of $r$, a reasonable heuristic is to compute a scalar score $o_{ij}$ for each edge $(i,j,\tau_{aij})$, and let the path’s score $o_{a}$ be the sum of the scores of all pairs of nodes that are connected by that path.
Following this intuition, for each node pair $(i,j)$, we concatenate $[\mathbf{y}_{i},\mathbf{y}_{j},\mathbf{e}_{ij},\tau_{(r+a)ij}]$ into a vector and apply an MLP to to compute the scalar $o_{ij}$. For each path $a\in\textup{SP}$, we sum these node-pair scores to get path scores: $o_{a}=\sum_{i,j\in a}o_{ij}$.
If $\alpha<1$, then the quality of $r$ still depends on the edges it adds to $\mathcal{E}_{\mathcal{R}}$, but also on its driving time, and whether it is a good choice to extend $r$ by $a$ depends generally on the state $\mathcal{S}$. So, we concatenate $[o_{a},\tau_{a},\mathbf{s}_{t}]$ and apply another MLP with 2 hidden layers to obtain a final score $\hat{o}_{a}$ for each candidate path. The extension policy is then the softmax of these values:
@@ A.4 Training Hyper-Parameters
Training was performed using the well-known Adam optimizer Kingma and Ba (2015). Table 7 gives the hyper-parameters used during training. In each epoch of training, we augment the training dataset by applying a set of random transformations. These include rescaling the node positions by a random factor uniformly sampled in the range $[0.4,1.6]$, rescaling the demand magnitudes by a random factor uniformly sampled in the range $[0.8,1.2]$, mirroring node positions about the $y$ axis with probability $0.5$, and revolving the node positions by a random angle in $[0,2\pi)$ about their geometric center.
We also randomly sample $\alpha$ separately for each training city in each epoch. Each time alpha is sampled, it has equal probability of being set to $0.0$, set to $1.0$, and sampled uniformly in the range $[0,1]$. This is to encourage the policy to learn both to handle intermediate cases, and the extreme cases of the passenger perspective and operator perspective.